OpenStudy (anonymous):

If water weighs 62 1/2 lbs per cubic foot, what is the weight of 14 1/4 cubic feet of water?

OpenStudy (mathmate):

Convert everything to an improper fraction and multiply. for example, 5 1/2 * 13 1/4 = \(\LARGE \frac{11}{2}\frac{53}{4}= \frac{583}{8}=72~7/8\)